Consider a woman puts Php 60000 in an account. After 3 years, she has a total amount of Php 70000. How long will it take to have Php 100000?


1
Expert's answer
2021-12-20T13:38:54-0500

"P_i=P(1+ \frac{r}{100})^n"

"70000=60000(1+ \frac{r}{100})^3"

"r=100(\sqrt[3]{70000/60000}-1)=5,27"

"100000=60000(1+\frac{5,27}{100})^n"

"n=log_{1.0527}(100000\/60000)=9.946"

Answer: It takes 9.946 year